OpenStudy (anonymous):

how do you do the ln(3x+5)=1

OpenStudy (aum):

If ln(a) = b, then a = e^b

OpenStudy (aum):

\[ \ln(3x+5) = 1 \implies 3x + 5 = e^1 = e \\ 3x + 5 = e \\ 3x = e - 5 \\ x = \frac 13(e - 5) \]